OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 10x + 11y = -26.5 20x + 5y = -27.5

OpenStudy (anonymous):

(10x + 11y = -26.5)mult by 2 20x+22y=-53 20x + 5y = -27.5 - (20x + 5y = -27.5) ------------------ 0+17y= -25.5 y=-25.5/17 y=-3/2 ans.. sub this to either of the eq or to 20x + 5y = -27.5 20x + 5(-3/2) = -27.5 20x - 15/2 = -27.5 20x=-20 x=-20/20 x=-1 ans (x,y)=(-1,-3/2) ,.....ans...

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw..good luck now